Shift summerie for OWL 11/11/1999

Arun Tripathi (arun@physics.ucla.edu)
Thu, 11 Nov 1999 05:11:59 -0800 (PST)


Shift Summary for owl shift, 11/11/99

On Shift: R. Barboas (00:00-3:40), A. Tripathi (3:40-7:00)

Hours of beam available at all intensities: 0.0
Hours of beam available at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:

========================================================================
Accesses during shift: None
========================================================================
Overview of shift: No beam, quiet.
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

Ricardo did the gas check. Spill signal comes back around
6:40. Tried starting a cosmic run, hangs at stage 3, did fresh
start, core dump from damp, another fresh start after removing
the core file and exiting from the window. In progress...

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift

Do a laser scan.

Grinder on espresso machine doesn't seem to be working. TOP PRIORITY!
========================================================================